Packaging Solutions Corporation manufactures and sells a wide variety of packaging products. Performance reports are prepared monthly for each department. The planning budget and flexible budget for the Production Department are based on the following formulas, where q is the number of labor-hours worked in a month:

Cost Formulas
Direct labor $16.50q
Indirect labor $4,700 + $1.30q
Utilities $5,200 + $0.60q
Supplies $1,700 + $0.20q
Equipment depreciation $18,700 + $2.60q
Factory rent $8,500
Property taxes $2,900
Factory administration $13,500 + $0.50q

The Production Department planned to work 4,400 labor-hours in March; however, it actually worked 4,200 labor-hours during the month. Its actual costs incurred in March are listed below:

Actual Cost Incurred in March
Direct labor $ 70,920
Indirect labor $ 9,600
Utilities $ 8,230
Supplies $ 2,790
Equipment depreciation $ 29,620
Factory rent $ 8,900
Property taxes $ 2,900
Factory administration $ 14,930

Required:

1. Prepare the Production Department’s planning budget for the month.

2. Prepare the Production Department’s flexible budget for the month.

3. Prepare the Production Department’s flexible budget performance report for March, including both the spending and activity variances.
Answer #1
1
Planning Budget
Budgeted labor hours 4400
Direct labor 72600 =4400*16.5
Indirect labor 10420 =4700+(4400*1.3)
Utilities 7840 =5200+(4400*0.6)
Supplies 2580 =1700+(4400*0.2)
Equipment Depreciation 30140 =18700+(4400*2.6)
Factory rent 8500
Property taxes 2900
Factory administration 15700 =13500+(4400*0.5)
Total expenses 150680
2
Flexible Budget
Actual labor hours 4200
Direct labor 69300 =4200*16.5
Indirect labor 10160 =4700+(4200*1.3)
Utilities 7720 =5200+(4200*0.6)
Supplies 2540 =1700+(4200*0.2)
Equipment Depreciation 29620 =18700+(4200*2.6)
Factory rent 8500
Property taxes 2900
Factory administration 15600 =13500+(4200*0.5)
Total expenses 146340
3
Actual Spending variance Flexible budget Activity variance Planning budget
Labor hours 4200 4200 4400
Direct labor 70920 1620 U 69300 3300 F 72600
Indirect labor 9600 560 F 10160 260 F 10420
Utilities 8230 510 U 7720 120 F 7840
Supplies 2790 250 U 2540 40 F 2580
Equipment Depreciation 29620 0 None 29620 520 F 30140
Factory rent 8900 400 U 8500 0 None 8500
Property taxes 2900 0 None 2900 0 None 2900
Factory administration 14930 670 F 15600 100 F 15700
Total expenses 147890 1550 U 146340 4340 F 150680